### v1

# Cost-Effective Text Clustering with Large Language Models

Abstract.
Text clustering aims to automatically partition a collection of text documents into distinct clusters based on linguistic features. In the literature, this task is usually framed as metric clustering based on text embeddings from pre-trained encoders or a graph clustering problem upon pairwise similarities from an oracle, e.g., a large ML model. Recently, large language models (LLMs) bring significant advancement in this field by offering contextualized text embeddings and highly accurate similarity scores, but meanwhile, present grand challenges to cope with substantial computational and/or financial overhead caused by numerous API-based queries or inference calls to the models.
In response, this paper proposes TECL, a cost-effective framework that taps into the feedback from LLMs for accurate text clustering within a limited budget of queries to LLMs. Under the hood, TECL adopts our EdgeLLM or TriangleLLM to construct must-link/cannot-link constraints for text pairs, and further leverages such constraints as supervision signals input to our weighted constrained clustering approach to generate clusters. Particularly, EdgeLLM (resp. TriangleLLM) enables the identification of informative text pairs (resp. triplets) for querying LLMs via well-thought-out greedy algorithms and accurate extraction of pairwise constraints through carefully-crafted prompting techniques. Our experiments on multiple benchmark datasets exhibit that TECL consistently and considerably outperforms existing solutions in unsupervised text clustering under the same query cost for LLMs.
Text Clustering, Constrained Clustering, Large Language Models
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Mathematics of computing Graph algorithmsccs: Information systems Clusteringccs: Information systems Language models
1. Introduction
Text clustering is a fundamental task in natural language processing (NLP), which seeks to group similar text documents into clusters. This task serves as a building block bolstering a wide range of practical applications, such as query refinement (Sadikov et al., 2010) in search engines, user intent detection (Zhang et al., 2021b; Rodriguez et al., 2024) in dialogue systems, social network analysis (Li and Li, 2024; Huang et al., 2022), topic discovery (Yin et al., 2011; Castellanos et al., 2017) and many others.
A canonical methodology for text clustering is to encode each text into an embedding vector, followed by applying metric clustering approaches, e.g., $k$-Means, hierarchical clustering, and DBSCAN, over these text representations to generate clusters (Aggarwal and Zhai, 2012). The effectiveness of such methods often hinges on the quality of text embeddings. The emergence of pre-trained language models (PLMs), e.g., BERT (Devlin, 2018) and its successors (Liu, 2019; Reimers, 2019; Su et al., 2023) has significantly advanced text clustering with contextual embeddings obtained via pre-training on a large corpus (Muennighoff et al., 2023). To date, large language models (LLMs) such as GPT4 (OpenAI, 2023), PaLM (Chowdhery et al., 2023), and LLaMA (Touvron et al., 2023) pre-trained on extensive text corpora and datasets at the expense of billions of parameters and vast computational resources achieve state-of-the-art text embeddings and have set new benchmarks in text clustering, as reported in (Petukhova et al., 2024; Keraghel et al., 2024). However, most of the powerful LLMs are deployed as paid online services, making the underlying text embeddings inaccessible or entailing computationally and financially expensive queries via APIs, particularly for a large text collection. Moreover, due to the restricted access to LLMs, text embeddings therefrom cannot be fine-tuned and customized for specific downstream tasks, rendering the rear-mounted clustering suboptimal.
With the advent of LLMs, another way for text clustering is to obtain similarity scores of text pairs from LLMs to define a graph clustering problem (Silwal et al., 2023). This methodology enables us to model the correlations between text instances with highly accurate semantic similarities/dissimilarities from LLMs, and hence, producing high-quality clustering results (Silwal et al., 2023; Zhang et al., 2023a). Despite the superb performance achieved, it requires pair-wise similarities between all text instances, leading to $O(n^{2})$ inference calls to the models, where $n$ is the total number of texts in the input text corpus, which is infeasible for clustering large-scale text collections in practice. Very recently, there are a number of works (Zhang et al., 2023a; Wang et al., 2023; Yang et al., 2024; Huang and He, 2024) exploring other ways of leveraging the feedback from instruction-tuned LLMs to guide text clustering. However, all of them still demand substantial queries to LLMs (typically with a token count of over 10×10$\times$10 × the corpus size) and/or rely on skeleton support from fine-tuning embedders such as BERT, Instructor, and E5, which is time-consuming.
Motivated by this, this paper proposes TECL (short for cost-effective TExt Clustering with LLMs), which aims to maximize the text clustering quality within a fixed budget (i.e., for token consumption) for querying LLMs. To be more specific, TECL is inspired by the constrained clustering paradigm (a.k.a. semi-supervised clustering) (Dinler and Tural, 2016), which additionally takes as input a number of constraints for pairs of data points, e.g., must-links or cannot-links, from human experts. As such, clusters can thus be obtained based on these constraints as supervision signals and cheap text embeddings from small PLMs, e.g., BERT or Sentence-BERT (Reimers, 2019). Given the human-level performance of LLMs in text comprehension, a simple and straightforward approach is to employ LLMs as the domain expert to create pair-wise pseudo constraints. However, this scheme poses two pressing challenges. First, it still remains unclear which text instances should be selected for querying and how to elicit high-caliber responses from LLMs. Secondly, the answers from LLMs might be noisy or random, engendering unreliable constraints that mislead the subsequent clustering.
To tackle the above problems, TECL includes two approaches EdgeLLM and TriangleLLM that both construct must-link and cannot-link constraints with LLMs in a cost-effective manner. EdgeLLM harnesses the prominent spanning edge centrality (Teixeira et al., 2013; Chandra et al., 1989) in graph theory to quantify the informativeness of text pairs (referred to as edges) and a well-thought-out greedy algorithm to achieve a linear-complexity identification of crucial edges. For each selected edge $(t_{a},t_{b})$, EdgeLLM elicits the relation, e.g., must-link (must be in the same cluster) or cannot-link (cannot belong to the same cluster), of $t_{a}$ and $t_{b}$ from LLMs with carefully-crafted prompts containing the task context and text contents. Based thereon, we further develop TriangleLLM that greedily selects crucial text triplets (referred to as triangles) for querying, which draws inspiration from the fact that humans represent an instance through comparing with others (Nosofsky, 2011). On top of that, TECL refines classic constrained clustering approaches, i.e., constrained spectral clustering (Wang et al., 2014) and constrained K-Means (Wang et al., 2014), as WCSC and WCKMeans, with our carefully designed penalty weights for pairwise constraints based on the pointwise mutual information (Church and Hanks, 1990) calculated from the text embeddings, so as to eliminate the side effects of misinformation introduced by LLMs.
Our extensive experiments comparing TECL against multiple strong baselines under the same query budget for LLMs on five benchmark text datasets demonstrate the consistent and superior performance of TECL in unsupervised text clustering tasks. In particular, on the well-known Tweet dataset, TECL is able to attain an accuracy of $73.1\%$ with a remarkable improvement of $7.29\%$ over the state-of-the-art at a monetary cost of \$0.0185 for calling GPT-4.0-Mini.
2. Related Work
In the sequel, we review existing studies germane to our work.
2.1. Text Clustering
Early methods for text clustering primarily rely on vector space models, e.g., bag-of-words and TF-IDF (Schütze et al., 2008), where documents are represented as high-dimensional sparse vectors and clustered using k-Means, DBSCAN, or other metric clustering algorithms. By transforming words into low-dimensional dense vectors based on their relationships in corpora, Word2Vec (Mikolov, 2013) and GloVe (Pennington et al., 2014) efficiently produce text embeddings encoding semantic and syntactic similarities, but still fail to capture contextual variations and polysemy. Transformer-based language models, e.g., BERT (Devlin, 2018) and its successors (Liu, 2019; Reimers, 2019; Su et al., 2023), mitigate the above issues by generating contextual embeddings via pre-training on a large corpus, which have set new benchmarks in text clustering.
Although the aforementioned embedding-based methods are simple and efficient, they fail to capture deeper semantic nuances in complex domains (Sun et al., 2019; Saha, 2023; Ravi and Kulkarni, 2023; Viswanathan et al., 2023; Muennighoff et al., 2023). A series of deep clustering approaches  (Zhou et al., 2022) are proposed to learn representations and cluster assignments jointly. They can be categorized into multiple groups, such as multi-stage (Tao et al., 2021; Huang et al., 2014), iterative (Yang et al., 2016; Caron et al., 2018; Van Gansbeke et al., 2020; Niu et al., 2022; Chang et al., 2017; Niu et al., 2020), generative (Dilokthanakul et al., 2016), and simultaneous methods (Xie et al., 2016; Zhang et al., 2021a; Hadifar et al., 2019), etc. While being effective, these approaches typically demand training on large datasets and careful tuning.
Recently, inspired by the powerful natural language capabilities of LLMs, several efforts have been made towards harnessing LLMs for enhanced text clustering. In Ref. (Keraghel et al., 2024; Petukhova et al., 2025), researchers have validated the superiority of LLMs over other text encoders when using their output text embeddings for clustering. Wang et al. (2023) propose to classify samples with explanations for clusters generated from LLMs, while De Raedt et al. (2023) collect descriptive labels using prototypical utterances to support in-context learning. ClusterLLM (Zhang et al., 2023a) achieves impressive clustering quality with sentence relations predicted by LLMs as the guide and finetuned embedders. Based on a subset of ground truth labels or expert feedback, Viswanathan et al. (2023) expand keyphrases and refine low-confidence cluster assignments to facilitate semi-supervised clustering, whose process pipeline is most similar to ours if reserving the LLM oracle and dropping other components. LLMEdgeRefine (Feng et al., 2024) creates superpoints to mitigate outliers and iteratively refines clusters using LLMs for improved semantic coherence. Yang et al. (2024) fine-tune embedding models based on the weakly-supervised contrastive learning loss and feedback from LLMs, followed by using LLMs to refine cluster assignments of low-confidence samples. Huang and He (2024) propose to transform the text clustering into a classification task by prompting LLMs to generate potential labels and assign the most appropriate label to each sample. Despite encouraging results reported, prior works strongly rely on extensive LLM queries, undergo costly fine-tuning, and/or require additional post-corrections, rendering them rather inefficient and impractical. Unlike previous works, our method achieves superior performance under identical conditions with a sufficiently low number of queries and budgets without fine-tuning or post-correction.
2.2. Constrained Clustering
To improve clustering performance, constrained clustering adopts side information such as pairwise must-link (ML) or cannot-link (CL) constraints. Classical approaches incorporate pairwise constraints into traditional clustering algorithms by refining their optimization frameworks. For example, constraint-based variants of K-means (Wagstaff et al., 2001) and Gaussian mixture models (Shental et al., 2003; Law et al., 2004, 2005) have been proposed to consider must-link and cannot-link relationships during clustering. A representative example is PCKMeans (Basu et al., 2004), which extends K-means by incorporating pairwise constraints into its objective function. However, these methods often struggle with robustness in high-dimensional embeddings or noisy datasets. Spectral methods represent another major branch, where constraints are encoded in graph structures (Kamvar et al., 2003; Kulis et al., 2005) or eigenspace manipulations (De Bie et al., 2004). These methods offer theoretically-grounded solutions but require a significant percentage of constraints and may overfit noise or inconsistencies in the constraint set. Among these, Constrained Spectral Clustering (CSC(Wang et al., 2014) stands out by integrating pairwise constraints into the spectral embedding through modifications to the graph Laplacian. Deep learning-based approaches address these limitations by leveraging neural networks to learn feature representations that respect constraints (Hsu and Kira, 2015; Chen, 2015). Methods such as SDEC (Ren et al., 2019) and constrained IDEC (Zhang et al., 2021c) refine latent spaces for clustering, but their discriminative nature limits their ability to model underlying data distributions. Generative extensions remain underexplored, with only a few studies (Luo et al., 2018) attempting to integrate constraints into probabilistic frameworks, often at the cost of computational efficiency. Despite progress, existing methods fail to fully exploit the structure of constraints in generative settings, leaving room for approaches that unify flexibility and interpretability in constrained clustering. In order to bridge this gap, we introduce weighting schemes based on PCKMeans and CSC and propose WCKMeans and WCSC including a pointwise mutual information weighting method, which elevates the constrained clustering from simply considering two types of constraints as equally crucial to specific weights for each constraint.
3. Problem Formulation
3.1. Notation and Terminology
Let $\mathcal{T}=\{t_{i}\}_{i=1}^{n}$ be an unlabeled corpus consisting of $n$ distinct text instances. We use $\Omega$ to denote the corpus size of $\mathcal{T}$, which is the total amount of text tokens in $\mathcal{T}$. The text embedding of text $t_{i}\in\mathcal{T}$ is denoted as $\overrightarrow{\mathbf{x}}_{i}\in\mathbb{R}^{d}$, which can be generated via any text encoder $f(t_{i})$, e.g., TF-IDF, BERT, and LLMs. We assume that each $\overrightarrow{\mathbf{x}}_{i}$ is $L_{2}$ normalized, i.e., $\|\overrightarrow{\mathbf{x}}_{i}\|_{2}=1$. Accordingly, we define the embedding-based semantic similarity (ESS) between any two text instances $t_{a},t_{b}\in\mathcal{T}$ as
(1) $s(t_{a},t_{b})=\textsf{cosine}(\overrightarrow{\mathbf{x}}_{a},\overrightarrow{\mathbf{x}}_{b})=\overrightarrow{\mathbf{x}}_{a}\cdot\overrightarrow{\mathbf{x}}_{b}.$
and the total similarities of a text $t_{a}\in\mathcal{T}$ with all others as
(2) $d(a)=\sum_{t_{i}\in\mathcal{T}}{s(t_{a},t_{i})}=\overrightarrow{\mathbf{x}}_{a}\sum_{t_{i}\in\mathcal{T}}{\overrightarrow{\mathbf{x}}_{i}}.$
We typically use $\{\mathcal{Y}_{1},\mathcal{Y}_{2},\ldots,\mathcal{Y}_{K}\}$ to represent a set of $K$ disjoint clusters and a cluster assignment matrix $\mathbf{Y}\in\mathbb{R}^{n\times K}$ to symbolize the text-cluster memberships. In particular, $\mathbf{Y}_{i,k}=1$ if $t_{i}\in\mathcal{Y}_{k}$ and 00 otherwise. A must-link (resp. cannot-link) refers to a pair of text instances $t_{i},t_{j}\in\mathcal{T}$ that are required to be in the same cluster (resp. distinct clusters). We use $[1,n]$ to denote the set of integers $\{1,2,\ldots,n\}$.
3.2. Problem Statement
Given text corpus $\mathcal{T}$, a query budget [^1] The total number of tokens for consumption. $Q$ to LLMs, and the desired (or ground-truth) number $K$ of clusters, our text clustering with constraints from LLMs (TCCL) task is to construct a must-link set (MLS) $\mathcal{M}$ and a cannot-link set (CLS) $\mathcal{C}$ based on the feedback from LLMs as supervision constraints within the query budget $Q$, such that the texts in $\mathcal{T}$ can be accurately partitioned into $K$ disjoint clusters $\{\mathcal{Y}_{1},\mathcal{Y}_{2},\ldots,\mathcal{Y}_{K}\}$, i.e., $\mathcal{Y}_{i}\cap\mathcal{Y}_{j}=\emptyset\ \forall{i,j\in[1,K]}$.
4. Synoptic Overview of TECL
Figure 1. An overview of TECL.
This section provides an overview of our proposed TECL framework for TCCL. As illustrated in Fig. 1, at a high level, TECL works in a cascaded pipeline, where the first step seeks to construct MLS $\mathcal{M}$ and CLS $\mathcal{C}$, i.e., must-links or cannot-links, via our text pair/triplet queries to LLMs, while the second step produces the cluster assignments for all text instances using the weighted constrained clustering algorithms that consider the importance weight of each constraint in $\mathcal{M}$ and $\mathcal{C}$.
Generating $\mathcal{M}$ and $\mathcal{C}$ via LLMs. More precisely, TECL relies on EdgeLLM or TriangleLLM to create MLS $\mathcal{M}$ and CLS $\mathcal{C}$, which build text pairs or triplets to query the LLMs to form pair-wise constraints. We refer to these two types of queries as edge and triangle queries henceforth, respectively. Denote by $s$ the average amount of tokens for a single text in $\mathcal{T}$. In EdgeLLM, we first generate
(3) $N={Q}/{(2s)}$
unique edges $\mathcal{S}$ from $\binom{n}{2}$ possible text pairs based on our edge centrality metric and greedy selection algorithm. Later, each of these edges $(t_{a},t_{b})\in\mathcal{S}$ will be sent to the LLM requesting an answer to whether they should belong to the same cluster or not. Accordingly, $|\mathcal{M}|+|\mathcal{C}|=N=Q/(2s)$ must-links or cannot-links can be extracted from these “Yes” or “No” answers, forming the MLS $\mathcal{M}$ and CLS $\mathcal{C}$.
Different from EdgeLLM, TriangleLLM selects
(4) $N_{\triangle}={Q}/{(3s)}$
distinct triangles $\mathcal{S}_{\triangle}$ from $\binom{n}{3}$ possible text triplets using a greedy approach. In particular, we ensure that any two triangles can have at most one common text so as to avoid duplicate pairs. TriangleLLM further obtains the relation (must-link or cannot-link) of every two texts in each $(t_{a},t_{b},t_{c})\in\mathcal{S}_{\triangle}$. Compared to edge queries, each triangle query yields three constraints. Hence, through TriangleLLM, we can generate $|\mathcal{M}|+|\mathcal{C}|=3\times N_{\triangle}=Q/s$ unique constraints in total, which is twice the amount as EdgeLLM, when given the same query budget $Q$.
Weighted Constrained Clustering. Classic constrained clustering approaches, e.g., Constrained Spectral Clustering (CSC(Wang et al., 2014) and PCKMeans (Basu et al., 2004), regard all must-links in $\mathcal{M}$ and cannot-links in $\mathcal{C}$ equally important. We extend them to their weighted variants, dubbed as WCSC and WCKMeans, by imposing personalized weights to various constraints based on their text embeddings. In doing so, we can offset the adverse impact of noisy must-links/cannot-links from LLMs.
Specifically, for any constraint on text pair $(t_{a},t_{b})$, we leverage the prominent pointwise mutual information (PMI) (Church and Hanks, 1990) to define the global important weight of $(t_{a},t_{b})$, which is formulated as follows:
(5) $\textsf{PMI}(t_{a},t_{b})=\log{\left(\frac{d(a)\cdot d(b)}{s(t_{a},t_{b})}\cdot\sum_{t_{x}\in\mathcal{T}}{\frac{1}{d(x)}}+1\right)}.$
Given a collection $\mathcal{E}$ of elements, the PMI of an element pair $(a,b)\in\mathcal{E}$ is defined as $\textsf{PMI}(a,b)=\log(\frac{\mathbb{P}(a,b)}{\mathbb{P}(a)\cdot\mathbb{P}(b)})$, where $\mathbb{P}(a)$ (resp. $\mathbb{P}(b)$) is the probability of observing $a$ (resp. $b$) in $\mathcal{E}$ and $\mathbb{P}(a,b)$ is the probability of observing pair $(a,b)$ in $\mathcal{E}$. The larger $\textsf{PMI}(a,b)$ is, the more likely $a$ and $b$ co-occur in $\mathcal{E}$. In Eq. (5), we let $\mathbb{P}(a)=\frac{1/d(a)}{\sum_{t_{x}\in\mathcal{T}}{\frac{1}{d(x)}}}$ (resp. $\mathbb{P}(b)=\frac{1/d(b)}{\sum_{t_{x}\in\mathcal{T}}{\frac{1}{d(x)}}}$) and $\mathbb{P}(a,b)=\frac{1/s(t_{a},t_{b})}{\sum_{t_{x}\in\mathcal{T}}{\frac{1}{d(x)}}}$, and add a shift value “1” to avoid negative output.
In WCSC, we first normalize the PMI values of pairs in $\mathcal{M}$ and $\mathcal{C}$ into a fixed range (e.g., $[0.5,1.5]$) and then construct the constraint matrix $\boldsymbol{\Theta}$ where $\boldsymbol{\Theta}_{a,b}=\textsf{PMI}(t_{a},t_{b})$, $-\textsf{PMI}(t_{a},t_{b})$ or 00 when $(t_{a},t_{b})$ is a must-link, cannot-link, or has no constraints, respectively. $\boldsymbol{\Theta}$ is later used to build the graph Laplacian matrix for subsequent spectral clustering.
In WCKMeans, aside from the standard within-cluster sum of squares objective function for K-means, it additionally incorporates the following penalty terms for constraints in the optimization:
(6) $\left(1+\sum_{(t_{a},t_{b})\in\mathcal{M}}w_{\mathcal{M}}(t_{a},t_{b})\right)+\left(1+\sum_{(t_{a},t_{b})\in\mathcal{C}}w_{\mathcal{C}}(t_{a},t_{b})\right),$
where $w_{\mathcal{M}}(t_{a},t_{b})$ (resp. $w_{\mathcal{C}}(t_{a},t_{b})$) stands for the penalty weight of must-link (resp. cannot-link) $(t_{a},t_{b})$. Particularly, we adopt $\textsf{PMI}(t_{a},t_{b})$ as the weight but further normalize it such that $w_{\mathcal{M}}\in[0.01,0.1]$ and $w_{\mathcal{C}}\in[0,0.01]$ to balance the impact of MLS and CLS.
In the succeeding Sections 5 and  6, we mainly focus on elaborating on the algorithmic details of our EdgeLLM and TriangleLLM methods for constructing MLS $\mathcal{M}$ and CLS $\mathcal{C}$.
5. The EdgeLLM Method
In this section, we first delineate our greedy algorithm in EdgeLLM for identifying the $N$ informative edges (Sections 5.1 and 5.2). Later, in Section 5.3, we introduce our carefully crafted prompts to LLMs and strategies to extract must-link and cannot-link constraints.
Figure 2. Illustration of the greedy edge/triangle selection.
5.1. Spanning Edge Centrality
To assess the informativeness and importance of each text pair in $\mathcal{T}\times\mathcal{T}$, we propose to capitalize on the spanning edge centrality (SEC) (a.k.a. effective resistance(Teixeira et al., 2013; Chandra et al., 1989) in graph theory as the selection criteria. Firstly, we build a full affinity graph $\mathcal{G}$ with $\mathcal{T}$ as the node set and $\mathcal{T}\times\mathcal{T}$ as the edge set, wherein the weight of each edge $(t_{a},t_{b})$ is defined by their ESS in Eq. (1). The weighted degree of any node $t_{a}\in\mathcal{T}$ is then $d(a)$ defined in Eq. (2).
Lemma 5.1 ((Lai et al., 2024)).
The SEC $\Phi(t_{a},t_{b})$ of each edge $(t_{a},t_{b})\in\mathcal{G}$ is bounded by
$\frac{1}{2}\left(\frac{1}{d(a)}+\frac{1}{d(b)}\right)\leq\Phi(t_{a},t_{b})\leq\frac{1}{1-\lambda_{2}}\left(\frac{1}{d(a)}+\frac{1}{d(b)}\right),$
where $\lambda_{2}<1$ stands for the second largest eigenvalue of the normalized adjacency matrix of $\mathcal{G}$.
In an unweighted graph $\mathcal{G}$, the SEC $\Phi(t_{a},t_{b})$ of an edge $(t_{a},t_{b})$ can be interpreted as the fraction of spanning trees of $\mathcal{G}$ that contains $(t_{a},t_{b})$. Intuitively, a large $\Phi(t_{a},t_{b})$ indicates that edge $(t_{a},t_{b})$ plays a critical role in the graph topology. However, the exact computation of $\Phi(t_{a},t_{b})$ is rather costly, and even most approximate algorithms still struggle to cope with medium-sized graphs (Spielman and Srivastava, 2008; Zhang et al., 2023b).
Lemma 5.1 states that $\Phi(t_{a},t_{b})$ of any edge $(t_{a},t_{b})$ is proportional to the sum of the inverses of two endpoints’ degrees, i.e.,
(7) $\Phi(t_{a},t_{b})\propto\frac{1}{d(a)}+\frac{1}{d(b)},$
which suggests us to use $\frac{1}{d(a)}+\frac{1}{d(b)}$ as a fast estimation of $\Phi(t_{a},t_{b})$ and has been shown accurate as validated in (Lai et al., 2024; Qiu et al., 2021). With slight abuse of notation, we let $\Phi(t_{a},t_{b})=\frac{1}{d(a)}+\frac{1}{d(b)}$ in the rest of this paper for tractability.
Input: Corpus $\mathcal{T}$ and the number $N$ of edges.
Output: An edge set $\mathcal{S}$.
1 Calculate the degree $t_{a}$ of each text $t_{a}\in\mathcal{T}$ by Eq. (2);
2 Sort $\mathcal{T}$ according to their degrees in the ascending order;
3 Initialize $p(t_{a})\leftarrow a+1$ for $1\leq a\leq\lceil\frac{\sqrt{8N+1}-1}{2}\rceil$;
4 Initialize a stack $\mathcal{S}=\{(t_{1},t_{2})\}$;
5 Update $p(t_{1})\leftarrow 3$;
6 while $|\mathcal{S}|<N$ do
7       $(t_{a},t_{b})\leftarrow\mathcal{S}.\textsf{top}()$;
8       ${b}^{-}\leftarrow p(t_{1}),\ {b}^{+}\leftarrow p(t_{a+1})$;
9       $\mathcal{B}\leftarrow\{(t_{1},t_{{b}^{-}}),(t_{a},t_{b+1}),(t_{a+1},t_{{b}^{+}})\}$;
10       $(t_{x},t_{y})\leftarrow\underset{(t_{a^{\prime}},t_{b^{\prime}})\in\mathcal{B}}{\operatorname{arg}\,\operatorname{max}}\;{\Phi(t_{a^{\prime}},t_{b^{\prime}})}$;
11       Push $(t_{x},t_{y})$ to $\mathcal{S}$;
12       $p(x)\leftarrow y+1$;
13
Algorithm 1 Greedy Edge Selection
5.2. Greedy Edge Selection
Let $\mathcal{A}=\{(t_{a},t_{b})\in\mathcal{T}\times\mathcal{T}|t_{a}\neq t_{b}\}$. Then, the edge selection task can be formally formulated as the following maximization problem:
(8) $\max_{\mathcal{S}\subset\mathcal{A},|\mathcal{S}|=N}\sum_{(t_{a},t_{b})\in\mathcal{S}}\Phi(t_{a},t_{b}),$
which seeks to extract a subset $\mathcal{S}\subseteq\mathcal{A}$ with $N$ distinct text pairs such that their total SEC over $\mathcal{G}$ is maximized.
A simple and straightforward way to get $\mathcal{S}$ is to calculate the SEC of each edge in $\mathcal{A}$ and pick the $N$-largest ones as $\mathcal{S}$. Such a brute-force approach entails a quadratic time complexity, i.e., $O(n^{2})$, to enumerate all text pairs in $\mathcal{A}$, which is impractical for large corpora. As a workaround, we resort to a greedy strategy for higher efficiency, whose pseudo-code is displayed in Algorithm 1.
Lemma 5.2.
$\underset{(t_{a},t_{b})\in\mathcal{S},\ a<b}{\operatorname{arg}\,\operatorname{max}}\;{a}\leq\lceil\frac{\sqrt{8N+1}-1}{2}\rceil$ when $|\mathcal{S}|=N$.
Proof.
Let $(t_{i},t_{j})=\underset{(t_{a},t_{b})\in\mathcal{S},\ a<b}{\operatorname{arg}\,\operatorname{max}}\;{a}$. In the extreme case, $j=i+1$ and only edges $(t_{a},t_{a+1}),(t_{a},t_{a+2}),\ldots,(t_{a},t_{j})$ are added into $\mathcal{S}$ for each $a\in[1,i-1]$ as they are surely larger than $(t_{i},t_{j})$. Put in another way, for each $t_{a}$ with $a\in[1,i]$, we can select $t_{b}$ with $a+1\leq b\leq j$, i.e., which corresponds to $j-a=i+1-a$ text instances. Thus, the total number of edges satisfy: $1+2+\ldots+i=\frac{(i+1)i}{2}=N$. Then, we can derive
$i^{2}+i-2N=0\Rightarrow i=\frac{\sqrt{8N+1}-1}{2},$
which finishes the proof. ∎
Specifically, Algorithm 1 begins by reordering the text instances in $\mathcal{T}$ according to their degrees defined in Eq. (2) in the ascending order, such that $\frac{1}{d(i)}$ of the $i$-th text $t_{i}$ is the $i$-th largest among all texts (Lines 1-2). Next, to facilitate the identification of the text $t_{b}$ that can yield the highest $\Phi(t_{a},t_{b})$ for each text $t_{a}$ ($b>a$), we maintain the index $p(t_{a})$ of the next suggested text for each possible $t_{a}$, which is initialized to be $a+1$ (Line 3). According to Lemma 5.2, the value of $a$ falls into the range of $[1,\lceil\frac{\sqrt{8N+1}-1}{2}\rceil]$. At Line 4, a stack $\mathcal{S}$ is additionally created to accommodate all the selected edges and $(t_{1},t_{2})$ is added to $\mathcal{S}$ as it has the globally largest SEC in $\mathcal{A}$. The index $p(t_{1})$ is then increased to $3$ since $t_{3}$ is the text that constitutes the new best edge with $t_{1}$ (Line 5). At a high level, the selection is to navigate between columns for the best edges in unvisited cells (in orange) as illustrated in Fig. 2(a).
Subsequently, Algorithm 1 starts an iterative process to greedily select the text pair $(t_{a},t_{b})$ ($b>a$) with the highest SEC to expand $\mathcal{S}$ (Lines 6-12). First, we get the latest added edge $(t_{a},t_{b})$ by picking the top element in stack $\mathcal{S}$ at Line 7. As such, the ideal edge $(t_{x},t_{y})$ to be selected in this round will be the one in $\mathcal{A}\setminus\mathcal{S}$ whose SEC is only second only to $\Phi(t_{a},t_{b})$. Simply, there are three possible choices for $(t_{x},t_{y})$, which satisfy $x=a$, $x>a$, or $x<a$, respectively. As per our greedy strategy, the first and second candidates are $(t_{a},t_{b+1})$ and $(t_{a+1},t_{b^{+}})$, where the index $b^{+}$ can be obtained via $p(t_{a+1})$. Intuitively, they are the two edges with the smallest SEC among all edges starting with $x=a$ and $x>a$, respectively. As for the case with $x<a$, Algorithm 1 greedily moves to the smallest value $x=1$ and selects $(t_{1},t_{b_{1}})$ ($b_{1}=p(t_{1})$) as the rest candidate (Line 8). We then form the set =$\mathcal{B}=\{(t_{1},t_{{b}^{-}}),(t_{a},t_{b+1}),(t_{a+1},t_{{b}^{+}})\}$ and find the one with the largest SEC as $(t_{x},t_{y})$, as in Lines 9-10. Next, at Lines 11-12, we push $(t_{x},t_{y})$ to $\mathcal{S}$ and update the index $p(t_{x})$ to $y+1$, i.e., the next best. Algorithm 1 repeats the above procedure until the size of $\mathcal{S}$ reaches $N$ (Line 6), and returns $\mathcal{S}$ as the final result.
Complexity Analysis. Since $\sum_{t_{i}\in\mathcal{T}}{\overrightarrow{\mathbf{x}}_{i}}$ in Eq. (2) can be precomputed and reused for the degree computations of all text instances in $\mathcal{T}$, Line 1 in Algorithm 1 takes $O(nd)$ time. The sorting at Line 2 can be done in $O(n\log{n})$ time using the quick sort algorithm. Note that Algorithm 1 will repeat Lines 6-12 for $N$ times, and each iteration only involves $O(1)$ operations. overall, the time complexity of Algorithm 1 can thus be bounded by $O(N+n\log{n}+nd)$.
Table 1. Example prompts for edge and triangle queries on CLINC.
Type Prompt
Edge Query Cluster CLINC docs by whether they are the same domain type. For each pair, respond with Yes or No without explanation.
- Domain #1: thank you ever so much for that!
- Domain #2: i want to eat something from turkey
Given this context, do domain #1 and domain #2 likely correspond to the same general domain type?
Triangle Query Cluster CLINC docs by whether they are the same domain type. For each triangle, respond with a, b, c, d, or e without explanation.
- Domain #1: i am interested in a new insurance plan
- Domain #2: go to the next song and play it
- Domain #3: how many days of vacation do i have left
Given this context, do domain #1, domain #2 and domain #3 likely correspond to the same general domain type?
(a) All are same domain.  (b) Only #1 and #2 are same domain.
(c) Only #1 and #3 are same domain.  (d) Only #2 and #3 are same domain.  (e) None.
Input: A min-heap $\mathcal{H}$ and new elements $\mathcal{E}$.
Output: The updated min-heap $\mathcal{H}$.
1 $h\leftarrow\mathcal{H}.\textsf{min}()$;
2 Push elements in $\mathcal{E}$ to $\mathcal{H}$;
3 if $\exists e\in\mathcal{E}\ s.t.\ e=h+1$ then
4       while $|\mathcal{H}|>1$ do
5             $\beta\leftarrow\mathcal{H}.\textsf{pop}()$;
6             if .$\mathcal{H}.\textsf{min}()\neq\beta+1$ then  Push $\beta$ to $\mathcal{H}$ and break ;
7
8
Algorithm 2 HeapAdd
Input: Two min-heaps $\mathcal{H}_{x}$ and $\mathcal{H}_{y}$.
Output: The minimum element $\gamma$.
1 if $\mathcal{H}_{y}.\textsf{min}()+1\notin\mathcal{H}_{x}$ then  $\gamma\leftarrow\mathcal{H}_{y}.\textsf{min}()+1$ ;
2else
3       for $\gamma\leftarrow\mathcal{H}_{y}.\textsf{min}()+2$ to $\mathcal{H}_{y}.\textsf{max}()+1$ do
4             if $\gamma\notin\mathcal{H}_{y}\cup\mathcal{H}_{x}$ then  break ;
5
6
Algorithm 3 GreedyScan
5.3. Edge Queries to LLMs
Given the edge set $\mathcal{S}$ selected by Algorithm 1, EdgeLLM starts to query the LLM (e.g., GPT-4) for the relation of the two text instances $t_{a},t_{b}$ in each edge $(t_{a},t_{b})$. Specifically, under the unsupervised setting, EdgeLLM first prompts the LLM with a context indicating the task being a clustering, the contents of text documents $t_{a},t_{b}$, and the reply format: “Yes” or “No” without providing explanation. Afterwards, the LLM is instructed to answer whether $t_{a}$ and $t_{b}$ should be in the same cluster. If the response is “Yes”, edge $(t_{a},t_{b})$ is recognized as a must-link and EdgeLLM adds it to MLS $\mathcal{M}$, otherwise edge $(t_{a},t_{b})$ is appended to CLS $\mathcal{C}$. Table 1 presents the detailed prompt template for edge queries.
6. The TriangleLLM Method
This section delineates the details of TriangleLLM in forming $N_{\triangle}$ triangles, querying the LLM, and extracting the MLS and CLS from the responses of the LLM.
Input: Corpus $\mathcal{T}$ and the number $N_{\triangle}$ of triangles.
Output: A triangle set $\mathcal{S}_{\triangle}$.
Lines 1-2 are the same as Lines 1-2 in Algorithm 1;
3 Initialize a min-heap $\mathcal{H}_{i}\leftarrow\{i\}\ \forall{i\in[1,n]}$;
4 Initialize a stack $\mathcal{S}_{\triangle}=\{(t_{1},t_{2},t_{3})\}$;
5 $\mathcal{H}_{1},\mathcal{H}_{2}\leftarrow\texttt{HeapAdd}(\mathcal{H}_{1},\{2,3\}),\ \texttt{HeapAdd}(\mathcal{H}_{2},\{3\})$;
6 while $|\mathcal{S}_{\triangle}|<N_{\triangle}$ do
7       $(t_{a},t_{b},t_{c})\leftarrow\mathcal{S}.\textsf{top}()$;
8       ${b}^{-},b^{\circ},{b}^{+}\leftarrow\mathcal{H}_{a-1}.\textsf{min}()+1,\mathcal{H}_{a}.\textsf{min}()+1,\mathcal{H}_{a+1}.\textsf{min}()+1$;
9       $c^{-}\leftarrow\texttt{GreedyScan}(\mathcal{H}_{a-1},\mathcal{H}_{b^{-}})$;
10       $c^{\circ}\leftarrow\texttt{GreedyScan}(\mathcal{H}_{a},\mathcal{H}_{b^{\circ}})$;
11       $c^{+}\leftarrow\texttt{GreedyScan}(\mathcal{H}_{a+1},\mathcal{H}_{b^{+}})$;
12       $\mathcal{B}\leftarrow\{(t_{a-1},t_{b^{-}},t_{c^{-}}),(t_{a},t_{b^{\circ}},t_{c^{\circ}}),(t_{a+1},t_{b^{+}},t_{c^{+}})\}$;
13       $(t_{x},t_{y},t_{z})\leftarrow\underset{(t_{a^{\prime}},t_{b^{\prime}},t_{c^{\prime}})\in\mathcal{B}}{\operatorname{arg}\,\operatorname{max}}\;{\Phi_{\triangle}(t_{a^{\prime}},t_{b^{\prime}},t_{c^{\prime}})}$;
14       Push $(t_{x},t_{y},t_{z})$ to $\mathcal{S}_{\triangle}$;
15       $\mathcal{H}_{x},\mathcal{H}_{y}\leftarrow\texttt{HeapAdd}(\mathcal{H}_{x},\{y,z\}),\ \texttt{HeapAdd}(\mathcal{H}_{y},\{z\})$;
16
Algorithm 4 Greedy Triangle Selection
6.1. Spanning Triangle Centrality Maximization
Inspired by the SEC in Section 5.1, we define the spanning triangle centrality (STC) of any triangle $(t_{a},t_{b},t_{c})$ in $\mathcal{G}$ as
(9) $\Phi_{\triangle}(t_{a},t_{b},t_{c})=\frac{\Phi(t_{a},t_{b})+\Phi(t_{a},t_{c})+\Phi(t_{b},t_{c})}{2}.$
Recall that $\Phi(t_{a},t_{b})$ signifies the fraction of spanning trees in $\mathcal{G}$ that include edge $(t_{a},t_{b})$. Analogously, $\Phi_{\triangle}(t_{a},t_{b},t_{c})$ can be roughly interpreted as the fraction of spanning trees of $\mathcal{G}$ where the triangle $(t_{a},t_{b},t_{c})$ appears. Using the of trick of estimating $\Phi(\cdot,\cdot)$ in Section 5.1, we reformulate $\Phi_{\triangle}(t_{a},t_{b},t_{c})$ as follows:
(10) $\Phi_{\triangle}(t_{a},t_{b},t_{c})=\frac{1}{d(a)}+\frac{1}{d(b)}+\frac{1}{d(c)}.$
Let $\mathcal{A}_{\triangle}$ be the subset of $\{(t_{a},t_{b},t_{c})\in\mathcal{T}\times\mathcal{T}\times\mathcal{T}|t_{a}\neq t_{b}\ \text{and}\ t_{b}\neq t_{c}\ \text{and}\ t_{a}\neq t_{c}\}$ such that any two distinct triangles $\triangle_{x},\triangle_{y}\in\mathcal{A}$ satisfy $|\triangle_{x}\cap\triangle_{y}|\leq 1$. Note that this requirement ensures that no duplicate text pairs are in all selected triangles. Similar in spirit to the objective for selecting edges in Eq. (8), the triangle selection is framed as the problem of selecting a subset $\mathcal{S}_{\triangle}$ of $N_{\triangle}$ triangles from $\mathcal{A}_{\triangle}$ such that their total STC is maximized:
(11) $\max_{\mathcal{S}_{\triangle}\subseteq\mathcal{A}_{\triangle},|\mathcal{S}_{\triangle}|=N_{\triangle}}\sum_{(t_{a},t_{b},t_{c})\in\mathcal{S}_{\triangle}}\Phi_{\triangle}(t_{a},t_{b},t_{c}).$
Enumerating all possible triangles in $\mathcal{A}_{\triangle}$ for the optimization of Eq. (11) yields a computation cost of $O(n^{3})$, which is prohibitively expensive. Instead, akin to EdgeLLM, we greedily select the triangle $(t_{a},t_{b},t_{c})$ ($a<b<c$) with the largest $\Phi_{\triangle}(t_{a},t_{b},t_{c})$. To facilitate this, for each text $t_{a}\in\mathcal{T}$, we need to maintain a set $\mathcal{H}_{a}$ comprising text instances that occur with $t_{a}$ in triangles already selected. Then, to find the new triangle $(t_{a},t_{b},t_{c})$, we can simply fetch the smallest element from $[a+1,n]\setminus\mathcal{H}_{a}$ as $t_{b}$ and then from $[b+1,n]\setminus(\mathcal{H}_{a}\cup\mathcal{H}_{b})$ as $t_{c}$, respectively. Unfortunately, this approach requires materializing $\mathcal{H}_{a}$ consisting of many elements and sorting $[a+1,n]\setminus\mathcal{H}_{a}$, which is both space-consuming and computationally inefficient.
Figure 3. The possible results of a triangle query.
6.2. Greedy Triangle Selection
To this end, we develop the Greedy Triangle Selection in Algorithm 4, which leverages a small-sized min-heap for each text $t_{a}\in\mathcal{T}$ to sidestep the redundant computations in searching for $t_{b}$ and $t_{c}$. As illustrated in Fig. 2(b), akin to Algorithm 1 for edge selection, the basic idea is to greedily determine the next move over the table containing candidate triangles.
As shown in Algorithm 4, we first reorder texts in $\mathcal{T}$ as Lines 1-2 in Algorithm 1. Then, a min-heap $\mathcal{H}_{i}=\{i\}$ is created for each text in $\mathcal{T}$ at Line 3, and a stack $\mathcal{S}_{\triangle}$ is initialized with adding the first triangle $(t_{1},t_{2},t_{3})$ to optimize Eq. (11) at Line 4. Accordingly, we update $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ using HeapAdd in Algorithm 2 (Line 5). Basically, rather than including the indices of all text instances occurring with $t_{a}$ in picked triangles in $\mathcal{S}_{\triangle}$, our idea is to merely keep the last “discontinuous” index $\beta$ and those larger than $\beta$. For example, instead of maintaining $\mathcal{H}_{2}=\{2,3,4,6,7\}$, we refine it as $\mathcal{H}_{2}=\{4,6,7\}$ via HeapAdd. In doing so, we can quickly identify the valid text $t_{b}$ for $t_{a}$ via $\mathcal{H}_{a}.\textsf{min}()+1$ for forming the triangle. In particular, Algorithm 2 iteratively pops out the minimum index $\beta$ from the min-heap $\mathcal{H}$ if the second smallest element is $\beta+1$, so as to the last “discontinuous” index.
Next, Algorithm 4 iteratively and greedily adds the new triangle to $\mathcal{S}_{\triangle}$ as follows. Based on the latest triangle $(t_{a},t_{b},t_{c})$ in $\mathcal{S}_{\triangle}$, we examine three candidates: $(t_{a-1},t_{b^{-}},t_{c^{-}})$, $(t_{a},t_{b^{\circ}},t_{c^{\circ}})$, and $(t_{a-1},t_{b^{+}},t_{c^{+}})$ (denoted as a set $\mathcal{B}$), one of whose STCs is expected to be second only to $\Phi_{\triangle}(t_{a},t_{b},t_{c})$ among those of $\mathcal{A}_{\triangle}\setminus\mathcal{S}_{\triangle}$. Equipped with $\mathcal{H}_{a-1}$, $\mathcal{H}_{a-1}$, and $\mathcal{H}_{a+1}$, we can easily get the best $t_{b^{-}}$, $t_{b^{\circ}}$, and $t_{b^{+}}$, respectively, as at Line 8.
As for $t_{c^{-}}$, $t_{c^{\circ}}$, and $t_{c^{+}}$, note that ${c^{-}}$, ${c^{\circ}}$, and ${c^{+}}$ cannot appear in $\mathcal{H}_{a-1}\cup\mathcal{H}_{b^{-}}$, $\mathcal{H}_{a}\cup\mathcal{H}_{b^{\circ}}$, and $\mathcal{H}_{a+1}\cup\mathcal{H}_{b^{+}}$, and should be greater than $\mathcal{H}_{b^{-}}.\textsf{min}()$, $\mathcal{H}_{b^{\circ}}.\textsf{min}()$, and $\mathcal{H}_{b^{+}}.\textsf{min}()$, respectively. To implement this, we develop GreedyScan in Algorithm 3, which takes input two min-heaps $\mathcal{H}_{x}$ and $\mathcal{H}_{y}$, and scans these two sets to find the minimum index $\gamma$ that is not in $\mathcal{H}_{x}\cup\mathcal{H}_{y}$ but $\gamma\geq\mathcal{H}_{y}.\textsf{min}()+1$. At Lines 9-11, we obtain ${c^{-}}$, ${c^{\circ}}$, and ${c^{+}}$ by invoking Algorithm 3.
Afterwards, we choose the triangle $(t_{x},t_{y},t_{z})$ with the maximal STC from $\mathcal{B}$ (Lines 12-13), and push it to $\mathcal{S}_{\triangle}$ at Line 14. At Line 15, the indices $\{y,z\}$ and $\{z\}$ are then inserted into the min-heaps $\mathcal{H}_{x}$ and $\mathcal{H}_{y}$ using HeapAdd, respectively. The above process (Lines 7-15) is repeated until $|\mathcal{S}_{\triangle}|=N_{\triangle}$, i.e., the $N_{\triangle}$ triangles are found.
Complexity Analysis. As per our analysis in Section 5.2, Lines 1-2 consume $O(n\log{n}+nd)$ time. The time cost of each iteration at Lines 7-15 is dependent on the overhead for invoking GreedyScan and HeapAdd functions. Assume that the maximum size of the heap $\mathcal{H}_{i}\ \forall{i\in[1,n]}$ is $h$, then each call to GreedyScan is bounded by $O(h)$ as it scans all elements in the heap in the worst case. In HeapAdd, each insertion to $\mathcal{H}$ takes $O(\log{h})$ time, and each deletion consumes $O(1)$ time. In the worst case, HeapAdd will pop out $h-1$ elements. Then, each invocation of HeapAdd requires up to $O(h+\log{h})$ time. Considering $N_{\Delta}$ iterations, the overall time complexity is therefore $O(n\log{n}+nd+N_{\Delta}h)$. Due to Lines 3-6, the empirical value of $h$ is rather small and thus can be regarded as a constant.
6.3. Triangle Queries to LLMs
Subsequently, TriangleLLM requests for relations of each text triplet $(t_{a},t_{b},t_{c})$ in $\mathcal{S}_{\triangle}$ returned by Algorithm 4 from the LLM using the following prompting approach. Akin to the edge queries in Section 5.3, the input prompts specify the task context and contents of text instances $t_{a},t_{b},t_{c}$, whereas the query is framed as a multiple-choice question (MCQ). In particular, the MCQ includes a set of five valid alternatives as depicted in the l.h.s. of Figure 3, wherein a line between two text instances indicates a must-link while the empty space represents a cannot-link. Note that the r.h.s. three results are invalid as they lead to contradictions. For instance, when $(t_{a},t_{b})$ and $(t_{a},t_{c})$ are must-links, it can be inferred that $(t_{b},t_{c})$ should be a must-link, which contradicts the fact that $(t_{b},t_{c})$ is a cannot-link. We instruct the LLM to respond with one of the five valid options (a)-(e) in Table 1 and add the corresponding must-links and cannot-links to $\mathcal{M}$ and $\mathcal{C}$ accordingly.
7. Experiments
This section experimentally evaluates the effectiveness of the proposed TECL framework and its constituent techniques (EdgeLLM, TriangleLLM, WCSC, and WCKMeans) in enhancing text clustering quality. All experiments are conducted on a Linux machine with an NVIDIA A100 GPU (80GB RAM), AMD EPYC 7513 CPU (2.6 GHz), and 1TB RAM.
Table 2. Dataset statistics.
Corpus #Texts ($n$) Corpus Size ($\Omega$) #Clusters ($K$) Task
BBC News 2,225 15,516 5 News
Tweet 2,472 27,883 89 Tweet
Bank77 3,080 38,773 77 Intent
Reddit 3,217 50,969 50 Topic
CLINC 4,500 39,970 10 Domain
7.1. Experimental Seup
Datasets. Table  2 presents the statistics of a wide range of benchmarks for various tasks that are widely used for experimental evaluation in previous works (Zhang et al., 2023a; Viswanathan et al., 2023; Zhang et al., 2021a). The corpus size $\Omega$ represents the total number of text tokens of the text instances in the respective five text corpora, which is also the token budget we want to control in the main experiments in Table 3 for fair comparison. The BBC News  (Keraghel et al., 2024) dataset contains 2,225 headlines, categorized into 5 topics, while the Tweet (Yin and Wang, 2016) dataset consists of 2,472 tweets labeled for relevance to 89 queries from the Text Retrieval Conference [^2] https://trec.nist.gov/data/microblog.html. Bank77  (Zhang et al., 2021b) includes 3,080 customer service utterances classified into 77 intent categories. The Reddit  (Muennighoff et al., 2023) dataset, with 3,217 posts, is used for topic discovery, and CLINC (Zhang et al., 2022) contains 4,500 requests labeled by 10 domains.
Baselines. We compare our proposed TECL against 10 baselines in text clustering. First, we include 7 embedding-based approaches that adopt 7 text embedders followed by K-Means++ or spectral clustering (Von Luxburg, 2007) for clustering, which include the classic TF-IDF, 4 popular PLMs: E5 (Wang et al., 2022), DistilBERT (Reimers, 2019), Sentence-BERT (Reimers, 2019), Instructor-Large (Su et al., 2023), and 2 LLMs: OpenAI-GPT (Brown et al., 2020) and LLaMA-2 (7B) (Touvron et al., 2023). Moreover, we consider three recent strong methods for text clustering: SCCL (Zhang et al., 2021a) utilizes contrastive learning to enhance cluster separation; ClusterLLM (Zhang et al., 2023a) leverages the feedback from an instruction-tuned LLM to fine-tune smaller embedding models and calibrate clustering granularity; PO-PCKMeans (Viswanathan et al., 2023) obtains pairwise constraints from LLMs with a subset of ground truth examples for semi-supervised clustering. For ClusterLLM, we also adopt its variant without fine-tuning for evaluation. The four versions of our TECL, i.e., EdgeLLM + WCKMeans, EdgeLLM + WCSC, TriangleLLM + WCKMeans, and TriangleLLM + WCSC, are included.
By default, we utilize Instructor-Large as the primary text embedder in SCCL, ClusterLLM, PO-PCKMeans, and TECL, due to its consistently superior performance. Additionally, we employ GPT-4o-Mini as the backbone LLM. For a fair comparison, we set the budget $Q$ for querying LLMs to the corpus size $\Omega$ on each dataset in TECL, ClusterLLM, PO-PCKMeans, and the embedding-based methods OpenAI-GPT and LLaMA-2 (7B). The parameters of all competitors are set as suggested in their respective papers. Following prior studies (Zhang et al., 2023a; Viswanathan et al., 2023; Zhang et al., 2021a; Feng et al., 2024), clustering accuracy (ACC) and normalized mutual information (NMI) are used as the clustering quality metrics.
Table 3. Clustering performance under the same query budget $Q=\Omega$ (best is highlighted in blue and runner-up underlined).
Method BBC News Tweet Bank77 Reddit CLINC
Embedding Clustering ACC NMI ACC NMI ACC NMI ACC NMI ACC NMI
TF-IDF K-means++ 24.88 1.84 55.19 77.62 36.35 58.83 12.87 18.66 25.38 14.09
SpectralClust 23.90 1.38 48.22 67.26 33.19 53.58 10.70 16.73 28.64 16.45
E5 K-means++ 71.66 53.03 60.90 85.91 58.74 77.09 38.88 48.75 59.60 56.25
SpectralClust 63.27 43.87 59.33 84.75 57.18 76.20 45.74 53.06 53.49 48.87
DistilBERT K-means++ 74.74 49.03 48.35 76.47 31.12 53.38 26.14 35.24 41.50 34.23
SpectralClust 71.42 41.82 51.08 77.62 39.71 60.45 30.40 38.44 35.47 27.79
Sentence-BERT K-means++ 79.02 56.88 62.74 86.42 60.77 78.84 42.31 50.02 50.32 48.98
SpectralClust 59.55 42.49 58.30 83.84 58.52 76.58 43.29 50.18 44.24 45.74
Instructor-Large K-means++ 86.53 71.51 65.34 87.74 64.57 82.01 53.62 62.62 55.94 57.41
SpectralClust 84.45 62.43 63.14 86.85 60.76 79.03 56.64 62.89 52.64 55.49
OpenAI-GPT K-means++ 88.56 73.65 65.42 87.64 64.20 81.74 51.06 60.10 55.31 56.05
SpectralClust 67.01 51.38 63.00 86.20 58.90 77.95 54.23 61.10 58.04 53.28
LLaMA-2 (7B) K-means++ 59.71 30.60 40.93 69.58 17.86 38.92 17.82 27.46 25.01 13.87
SpectralClust 49.05 24.11 52.29 76.27 27.60 49.33 24.43 34.67 27.08 18.46
SCCL (Zhang et al., 2021a) 83.60 62.00 36.50 68.50 36.00 61.30 28.70 36.00 46.60 44.00
ClusterLLM (Zhang et al., 2023a) 88.40 69.73 65.81 88.61 68.27 82.45 54.80 62.47 51.90 55.04
ClusterLLM w/o finetune (Zhang et al., 2023a) 89.07 70.97 64.59 87.78 65.14 82.33 54.75 62.87 58.66 58.95
PO-PCKMeans (Viswanathan et al., 2023) 89.53 72.16 65.58 88.20 67.79 81.92 55.10 62.51 57.80 56.32
TECL (EdgeLLM + WCKMeans) 90.47 73.96 71.16 88.90 65.29 82.52 57.07 63.80 62.31 61.11
TECL (EdgeLLM + WCSC) 86.88 69.49 70.91 88.51 67.99 82.55 57.01 63.21 56.20 57.77
TECL (TriangleLLM + WCKMeans) 90.65 74.25 72.25 88.60 66.72 83.07 59.28 64.98 65.16 62.09
TECL (TriangleLLM + WCSC) 91.01 75.22 73.10 89.08 69.03 83.20 57.78 64.25 57.91 60.02
Table 4. Ablation study on TECL.
Variant BBC News Tweet
ACC NMI ACC NMI
w/o WCSC 90.43 73.9 65.33 87.84
w/o TriangleLLM 88.72 70.25 67.27 87.44
TECL (TriangleLLM + WCSC) 91.01 75.22 73.10 89.08
TECLClusterLLMPO-PCKmeans
${\Omega}/{10}$${\Omega}/{5}$${\Omega}/{2}$$\Omega$$2\Omega$0.870.880.890.900.910.92query budgetACC
((a)) BBC News
${\Omega}/{10}$${\Omega}/{5}$${\Omega}/{2}$$\Omega$$2\Omega$0.620.640.660.680.700.720.74query budgetACC
((b)) Tweet
Figure 4. Clustering performance when varying $Q$
7.2. Text Clustering Performance
Table 3 reports the ACC and NMI scores achieved by all the competitors and our TECL on the five datasets. The best and second-best results are highlighted in blue, and darker shades indicate better clustering. The best baselines are underlined. Specifically, we can make the following observations. First, TECL consistently and considerably outperforms all the baselines over all datasets, even including the semi-supervised method PO-PCKMeans. For example, on Tweet and CLINC datasets, TECL can achieve a significant improvement of $7.29\%$ and $5.56\%$ in terms of ACC compared to the best baseline results, respectively. The remarkable gains validate the effectiveness of our proposed framework, which leverages greedy edge/triangle selection algorithms and LLMs to create informative pair constraints, followed by weighted constrained clustering. Amid the four versions of TECL, TriangleLLM is always superior to EdgeLLM since the former yields more pair constraints compared to the latter under the same query budget $Q$, as pinpointed in Section 4. On BBC News, Tweet, and Bank77 datasets, it can be seen that WCSC produces better clusters than WCKMeans, while the latter outperforms the former on Reddit and CLINC. Unless otherwise specified, we refer to TECL with TriangleLLM, WCSC as TECL for simplicity in the remaining experiments.
Another key observation we can make from Table 3 is that the performance of most embedding-based approaches is subpar, except Instructor-Large and OpenAI-GPT, and Instructor-Large is even superior or comparable to OpenAI-GPT on multiple datasets, which reveals that small PLMs can also output high-quality text embeddings for clustering. However, both of them are still inferior to other LLM-based approaches, including ClusterLLM, PO-PCKMeans, and our TECL. The results manifest that the direct adoption of the embeddings from LLMs for text clustering is less than satisfactory, whereas using the explicit correlations between text instances from LLMs shows higher effectiveness, which is consistent with the observation in other studies (Silwal et al., 2023; Zhang et al., 2023a).
Table 5. TECL with various selection strategies.
Dataset Strategy ACC NMI
BBC News Max (Ours) 91.01 75.22
Min 67.42 64.70
Random 88.72 70.25
Tweet Max (Ours) 73.10 89.08
Min 66.63 87.12
Random 67.27 87.44
7.3. Performance when Varying Query Budgets
In this set of experiments, we further study the text clustering performance of TECL and other LLM-based baselines ClusterLLM and PO-PCKMeans when the query budget $Q$ is varied. Fig. 4 depicts their ACC results on BBC News and Tweet when varying $Q$ in $\{\frac{\Omega}{10},\frac{\Omega}{5},\frac{\Omega}{2},\Omega,2\Omega\}$. It can be observed that ClusterLLM is unable to leverage the feedback from LLMs effectively, as its performance slightly decreases when $Q$ increases. A similar issue can be observed in PO-PCKMeans, whose ACC scores undergo a remarkable decline when $Q$ is increased from $\Omega$ (resp. $\frac{\Omega}{5}$) to $2\Omega$ (resp. $\frac{\Omega}{2}$) on BBC News (resp. Tweet). In sum, both baselines fail to obtain notable improvements when expanding the budget for querying LLMs.
WCSC (Ours)CSCPCKMeansCOPKMeansCCPCKMeansMPCKMeans
0.700.740.780.820.860.900.94ACC
((a)) BBC
0.580.620.660.70.74ACC
((b)) Tweet
Figure 5. TECL with various constrained clustering methods
By contrast, a consistent and conspicuous performance rise can be observed for TECL on both tested datasets, and the ACC scores are always considerably higher than those by ClusterLLM and PO-PCKMeans under the same query budget. In particular, when the query budget $Q=\frac{\Omega}{10}$, TECL achieves the clustering performance that is even superior to those attained by ClusterLLM and PO-PCKMeans with $Q=2\Omega$ on both datasets, indicating the high query effectiveness of TECL.
Table 6. Various weighting schemes for constraints.
Method Weight BBC News Tweet
ACC NMI Acc NMI
WCSC None 90.43 73.90 65.33 87.84
ESS 72.40 65.31 66.55 87.22
SESS 71.37 63.35 66.34 87.17
LESS 71.15 64.36 65.33 86.85
SEC 69.48 63.82 67.27 87.63
SSEC 70.07 65.21 67.72 87.28
IPMI 73.12 65.91 68.89 88.01
PMI (Ours) 91.01 75.22 73.10 89.08
WCKmeans None 89.80 72.38 65.66 87.84
ESS 90.11 73.41 64.00 87.52
SESS 89.98 73.08 63.63 87.41
LESS 90.11 73.33 64.56 87.00
SEC 88.85 70.83 63.71 86.90
SSEC 89.03 70.93 63.63 87.04
IPMI 90.47 74.03 66.63 87.45
PMI (Ours) 90.65 74.25 72.25 88.60
7.4. Ablation Study and Component Analysis
This section presents ablation studies and the efficacy of each ingredient in TECL.
Ablation Study. We conduct an ablation study to explore the effects of WCSC proposed in Section  4 and TriangleLLM in Section  6.
As shown in Table 4, the removal of WCSC results in a decrease in performance, especially on Tweet (ACC falls from 73.10% to 65.33%). This indicates that the weighting of constraints helps in managing noisy or less informative pairs. Additionally, the absence of TriangleLLM also causes a drop in performance, especially on BBC News (ACC declines from 91.01% to 88.72%), highlighting the importance of selecting high-quality constraints. These findings suggest that both components play a significant role.
Edge/Triangle Selection Strategies. We study three strategies in TriangleLLM to select triangles for querying LLMs, where the Max (resp. Min) is to maximize (resp. minimize) the SEC scores, whilst the Random generates triangles randomly. Table 5 shows the Max strategy as implemented by Algorithm 4 achieves the best performance (e.g., 91.01% ACC on BBC News). In comparison, the Min and Random strategies show significantly inferior performance due to selecting low-SEC or random triangles. The results validate the effectiveness of our greedy selection scheme proposed in Section 6.
Constrained Clustering Methods. Fig. 5 presents the clustering performance of TECL with various constrained clustering methods, including WCSC (ours), CSC (Wang et al., 2014), PCKMeans (Basu et al., 2004), COPKMeans (Bibi et al., 2023), CCPCKMeans (Rujeerapaiboon et al., 2019), and MPCKMeans (Bilenko et al., 2004) on BBC News and Tweet. Our WCSC that assign personalized weights to constraints achieves the highest accuracy (91.01% on BBC and 73.10% on Tweet), indicating its superiority to exploit pairwise constraints.
Weighting Schemes for Constraints. In Table 6, we report the results of TECL when using various weighting schemes in our WCSC and WCKmeans on BBC News and Tweet. The additional weighting schemes for comparison are listed in the following table:
ESS SESS LESS
$s(t_{a},t_{b})$ $\sqrt{s(t_{a},t_{b})}$ $\log{(s(t_{a},t_{b})+1)}$
SEC SSEC IPMI
$\frac{1}{d(a)}+\frac{1}{d(b)}$ $\sqrt{\frac{1}{d(a)}+\frac{1}{d(b)}}$ $\log{\left(\frac{s(t_{a},t_{b})\cdot\sum_{t_{x}\in\mathcal{T}}{d(x)}}{{d(a)}\cdot{d(b)}}+1\right)}$
Table 6 shows that improper weights for constraints lead to a significant adverse impact on the performance in both WCSC and WCKmeans. Amid all weighting schemes, the PMI defined in Section 4 consistently achieves the best results. For example, on Tweet, using PMI in WCSC improves over its unweighted version (i.e., None) by a large margin of $7.77\%$, while the best competitor IPMI can only attain an improvement of $3.56\%$.
Table 7. Clustering performance with various text encoders.
Encoder Method BBC News Tweet
ACC NMI ACC NMI
Sentence-BERT SCCL 82.00 58.33 53.60 72.80
ClusterLLM 80.08 54.13 57.47 83.74
PO-PCKMeans 82.70 57.43 61.39 86.12
TECL 84.45 61.65 78.48 89.30
Instructor-Large SCCL 83.60 62.00 36.50 68.50
ClusterLLM 88.40 69.73 65.81 88.61
PO-PCKMeans 89.53 72.16 65.58 88.20
TECL 91.01 75.22 73.10 89.08
Various Text Encoders. Table 7 conveys the clustering performance of TECL and recent baselines with two strong text encoders. The results show the superior effectiveness of TECL: (1) TECL achieves state-of-the-art performance in various encoders, achieving 91. 01% ACC on BBC News with Instructor-Large and 78.48% ACC on Tweet with Sentence-BERT, surpassing the second-best baselines. (2) Under the same encoder, TECL consistently outperforms competitors: with Sentence-BERT, TECL exceeds the second-best method (PO-PCKMeans) with a gain of 1.75% in ACC on BBC News; with Instructor-Large, TECL surpasses ClusterLLM by an improvement of 7.29% in ACC on Tweet. These gains, coupled with superior NMI scores, validate TECL’s ability to leverage pretrained embeddings and cherry-pick constraints without fine-tuning.
GPT-4oGPT-4o-Mini
Gemini-1.5-ProClaude-3.5-Sonnet
0.80.820.840.860.880.90.92Acc
((a)) BBC News
0.680.690.70.710.720.730.74Acc
((b)) Tweet
Figure 6. The performance of TECL with various LLMs
Queried LLMs. Fig. 6 illustrates the clustering results of TECL when querying 4 popular LLMs: GPT-4o, GPT-4o-Mini, Gemini-1.5-Pro, and Claude-3.5-Sonnet. Key observations include: (1) Minimal differences in the adoption of various LLMs. (2) LLMs influence answer accuracy, particularly regarding constraints, without affecting query triangle selection; for example, GPT-4o enhances constraint accuracy compared to GPT-4o-Mini, resulting in increased overall accuracy. (3) The advantages of more expensive LLMs do not warrant their costs, underscoring the necessity for more cost-effective models capable of managing complex tasks like triangle queries.
8. Conclusion
This paper presents TECL, a framework that cost-effectively leverages LLMs to produce accurate pair-wise constraints for subsequent constrained text clustering. In particular, we develop two greedy algorithms EdgeLLM and TriangleLLM to carefully select text pairs and triplets for querying LLMs so as to obtain the most informative constraints for clustering, and further ameliorate standard constrained clustering algorithms with personalized weights for various pair constraints. Our extensive experiments over benchmark datasets exhibit the high superiority of our proposed techniques in enhancing text clustering quality and reducing the sheer amount of queries needed for LLMs.